2 12 12 12 17 13 0 5 13 10 14 11 6 18 6 17...

2

12

12

12

17

13

0

5

13

10

14

11

6

18

6

17

6

5

4

5

What is the expected number of runs?

A)

5.45

B)

8.9

C)

9.9

D)

10.9
Answer #1

2,12,12,12,17,13,0,5,13,10,14,11,6,18,6,17,6,5,4,5

expected number =

x/n = 198/20

= 9.9

expected number is 9.9